# Investigating and Defending Shortcut Learning in Personalized Diffusion Models

### v2

@@ 4.2 Systematic Defending with Input Purification, Decouple Learning, and Guided Sampling
Motivated by our empirical observation, we propose two directions to enhance the robustness when learning personalized diffusion on a potentially corrupted dataset: i) weakening those spurious paths toward noisy prediction; ii) enhancing those toward prior data $\bar{\mathbf{x}}_{0}$ and true instance data $\mathbf{x}_{0}$. For the first direction, we leverage the off-the-shelf super-resolution and image restoration models to improve the data quality and remove the noise $\bm{\delta}$. For the second part, we introduce a novel noise-aware contrastive learning with negative tokens to make the model focus on those concept-related features and avoid being trapped in the local minimum of learning perturbation. Furthermore, during the sampling, we guide to generate high-quality images with simple negative prompting. We present our overall framework in Figure 4 and describe each of the propose modules in details as follow.
@@ 5.1 Experimental Setup
Table 1: Results of different purification methods under different protective perturbations. The best performances are in bold, and second runners are shaded in gray.
@@
Clean Data 0.27 0.21 0.27 0.21 0.27 0.21 0.27 0.21 0.27 0.21 0.27 0.21 0.27 0.21
Perturbed Data -0.02 -0.61 -0.04 -0.62 -0.04 -0.60 0.10 -0.60 0.01 -0.29 0.17 -0.13 0.25 -0.19
Gaussian F. 0.14 -0.32 0.15 -0.34 0.14 -0.33 0.15 -0.44 0.15 -0.35 0.14 -0.28 0.12 -0.24
JPEG 0.11 -0.46 0.14 -0.43 0.13 -0.43 0.06 -0.47 0.19 -0.01 0.23 0.15 0.16 0.24
TVM 0.03 -0.45 0.05 -0.47 0.09 -0.47 0.06 -0.50 0.07 -0.47 0.04 -0.50 -0.09 -0.48
PixelDiffPure -0.04 -0.37 0.01 -0.42 -0.05 -0.49 -0.04 -0.44 0.00 -0.42 0.04 -0.44 -0.09 -0.57
LatentDiffPure-$\emptyset$ 0.06 -0.45 -0.03 -0.50 0.07 -0.42 -0.07 -0.28 -0.13 -0.18 0.04 -0.28 -0.09 -0.21
LatentDiffPure 0.02 -0.36 -0.02 -0.52 0.02 -0.41 -0.05 -0.26 -0.09 -0.18 -0.02 -0.26 -0.13 -0.22
DDSPure 0.17 -0.23 0.20 -0.28 0.23 -0.26 0.15 -0.48 0.16 -0.30 0.17 -0.25 0.13 -0.25
GrIDPure 0.17 -0.11 0.18 -0.09 0.20 -0.16 0.16 -0.17 0.19 -0.07 0.19 -0.10 0.28 -0.08
IMPRESS 0.12 -0.55 0.11 -0.54 0.08 -0.56 0.11 -0.60 0.06 -0.35 0.20 -0.10 0.28 -0.45
Ours 0.27 0.45 0.27 0.34 0.28 0.35 0.29 0.37 0.16 0.49 0.19 0.47 0.23 0.48
@@ 5.4 Ablation Study
6 Conclusion, Discussion, and Future Works
@@ C.2 Protective Perturbation Methods
We test a wide range of protective perturbation approaches, including those that craft noise against fixed LDMs by exploiting the out-of-distribution adversarial vulnerability of DNNs [7, 57, 28, 5, 6], and those that jointly and alternatively learn the noise generator and perturbation [4, 8, 29], which show more protection capacity for the LDM fine-tuning settings [22, 23].
@@
Ensemble-based ASPL (EASPL). Since the model trainer’s pre-trained text-to-image generator is often unknown, an improved approach is to use an ensemble [61, 62] of surrogate models finetuned from different pre-trained generators, which can lead to better transferability. We implement this approach with three surrogates, including, SD v1-4, SD v1-5 and SD v2-1-base. Besides, we follow the practice of a single model at a time in an interleaving manner to produce optimal perturbed data due to GPU memory constraints.
MetaCloak. Despite the effectiveness of perturbation crafted from noise-surrogate joint learning, studies find that these approaches lack robustness against simple data transformations such as minor Gaussian filtering. To address this issue, MetaCloak [8] solves the underlying bi-level poisoning problem using a meta-learning framework with an additional transformation sampling process to craft transferable and robust perturbations. A pool of steps-staggered surrogate diffusion models is employed to create model-agnostic perturbations. By incorporating an additional transformation process and a denoising-error maximization loss, MetaCloak causes severe performance degradation in a personalized generation.

### v3

@@ 4.2 Systematic Defending with Input Purification, Decouple Learning, and Guided Sampling
Motivated by our empirical observation, we propose two directions to enhance the robustness when learning personalized diffusion on a potentially corrupted dataset: i) weakening those spurious paths toward noisy prediction; ii) enhancing those toward prior data $\bar{\mathbf{x}}_{0}$ and true instance data $\mathbf{x}_{0}$. For the first direction, we leverage the off-the-shelf super-resolution and image restoration models to improve the data quality and remove the noise $\bm{\delta}$. For the second part, we introduce a novel noise-aware contrastive learning with negative tokens to make the model focus on those concept-related features and avoid being trapped in the local minimum of learning perturbation. Furthermore, during the sampling, we guide to generate high-quality images with simple negative prompting. We present our overall framework in Figure 4 and Alg. 1 and describe each of the proposed modules in detail as follows.
@@ 5.1 Experimental Setup
Table 1: Results of different purification methods under different protective perturbations. The best performances are in bold, and second runners are shaded in gray. * denotes significant improvement that passes the Wilcoxon signed-rank significance test with $p\leq 0.01$.
@@
Clean 0.27 $\pm$ 0.10 0.21 $\pm$ 0.11 0.27 $\pm$ 0.10 0.21 $\pm$ 0.10 0.27 $\pm$ 0.10 0.21 $\pm$ 0.10 0.27 $\pm$ 0.10 0.21 $\pm$ 0.12 0.27 $\pm$ 0.10 0.21 $\pm$ 0.11 0.27 $\pm$ 0.10 0.21 $\pm$ 0.11 0.27 $\pm$ 0.10 0.21 $\pm$ 0.11
Perturbed -0.02 $\pm$ 0.09 -0.61 $\pm$ 0.07 -0.04 $\pm$ 0.08 -0.62 $\pm$ 0.12 -0.04 $\pm$ 0.08 -0.60 $\pm$ 0.12 0.10 $\pm$ 0.10 -0.60 $\pm$ 0.08 0.01 $\pm$ 0.09 -0.29 $\pm$ 0.09 0.17 $\pm$ 0.09 -0.13 $\pm$ 0.08 0.25 $\pm$ 0.10 -0.19 $\pm$ 0.10
Gaussian F. 0.14 $\pm$ 0.10 -0.32 $\pm$ 0.08 0.15 $\pm$ 0.10 -0.34 $\pm$ 0.06 0.14 $\pm$ 0.10 -0.33 $\pm$ 0.07 0.15 $\pm$ 0.09 -0.44 $\pm$ 0.04 0.15 $\pm$ 0.09 -0.35 $\pm$ 0.07 0.14 $\pm$ 0.10 -0.28 $\pm$ 0.08 0.12 $\pm$ 0.09 -0.24 $\pm$ 0.07
JPEG 0.11 $\pm$ 0.10 -0.46 $\pm$ 0.10 0.14 $\pm$ 0.11 -0.43 $\pm$ 0.07 0.13 $\pm$ 0.11 -0.43 $\pm$ 0.07 0.06 $\pm$ 0.11 -0.47 $\pm$ 0.12 0.01 $\pm$ 0.11 -0.01 $\pm$ 0.11 0.12 $\pm$ 0.11 0.15 $\pm$ 0.11 0.16 $\pm$ 0.10 0.24 $\pm$ 0.14
TVM 0.03 $\pm$ 0.08 -0.45 $\pm$ 0.06 0.05 $\pm$ 0.09 -0.47 $\pm$ 0.07 0.09 $\pm$ 0.09 -0.47 $\pm$ 0.06 0.06 $\pm$ 0.09 -0.50 $\pm$ 0.04 0.07 $\pm$ 0.09 -0.47 $\pm$ 0.06 0.04 $\pm$ 0.09 -0.50 $\pm$ 0.06 -0.09 $\pm$ 0.08 -0.48 $\pm$ 0.06
PixelDiffPure -0.04 $\pm$ 0.08 -0.37 $\pm$ 0.05 0.01 $\pm$ 0.08 -0.42 $\pm$ 0.06 -0.05 $\pm$ 0.09 -0.49 $\pm$ 0.05 -0.04 $\pm$ 0.08 -0.44 $\pm$ 0.06 0.00 $\pm$ 0.09 -0.42 $\pm$ 0.06 0.04 $\pm$ 0.09 -0.44 $\pm$ 0.04 -0.09 $\pm$ 0.08 -0.57 $\pm$ 0.05
L.DiffPure-$\emptyset$ 0.06 $\pm$ 0.09 -0.45 $\pm$ 0.08 -0.03 $\pm$ 0.09 -0.50 $\pm$ 0.06 0.07 $\pm$ 0.10 -0.42 $\pm$ 0.08 -0.07 $\pm$ 0.08 -0.28 $\pm$ 0.08 -0.13 $\pm$ 0.08 -0.18 $\pm$ 0.11 0.04 $\pm$ 0.08 -0.28 $\pm$ 0.08 -0.09 $\pm$ 0.08 -0.21 $\pm$ 0.08
L.DiffPure 0.02 $\pm$ 0.08 -0.36 $\pm$ 0.09 -0.02 $\pm$ 0.09 -0.52 $\pm$ 0.06 0.02 $\pm$ 0.09 -0.41 $\pm$ 0.09 -0.05 $\pm$ 0.08 -0.26 $\pm$ 0.06 -0.09 $\pm$ 0.08 -0.18 $\pm$ 0.10 -0.02 $\pm$ 0.08 -0.26 $\pm$ 0.06 -0.13 $\pm$ 0.08 -0.22 $\pm$ 0.08
DDSPure 0.17 $\pm$ 0.11 -0.23 $\pm$ 0.07 0.20 $\pm$ 0.10 -0.28 $\pm$ 0.06 0.23 $\pm$ 0.09 -0.26 $\pm$ 0.06 0.16 $\pm$ 0.10 -0.48 $\pm$ 0.07 0.13 $\pm$ 0.10 -0.30 $\pm$ 0.07 0.17 $\pm$ 0.10 -0.25 $\pm$ 0.06 0.13 $\pm$ 0.10 -0.25 $\pm$ 0.07
GrIDPure 0.17 $\pm$ 0.10 -0.11 $\pm$ 0.07 0.18 $\pm$ 0.10 -0.09 $\pm$ 0.07 0.20 $\pm$ 0.10 -0.16 $\pm$ 0.07 0.12 $\pm$ 0.09 -0.17 $\pm$ 0.07 0.02 $\pm$ 0.09 -0.07 $\pm$ 0.08 0.13 $\pm$ 0.10 -0.10 $\pm$ 0.08 0.28 $\pm$ 0.10 -0.08 $\pm$ 0.08
IMPRESS 0.12 $\pm$ 0.10 -0.55 $\pm$ 0.07 0.11 $\pm$ 0.10 -0.54 $\pm$ 0.09 0.08 $\pm$ 0.10 -0.56 $\pm$ 0.04 0.11 $\pm$ 0.10 -0.60 $\pm$ 0.09 0.06 $\pm$ 0.10 -0.35 $\pm$ 0.09 0.11 $\pm$ 0.10 -0.10 $\pm$ 0.08 0.28 $\pm$ 0.10 -0.45 $\pm$ 0.07
Ours 0.27 $\pm$ 0.08 0.45 $\pm$ 0.06 0.27 $\pm$ 0.08 0.34 $\pm$ 0.04 0.28 $\pm$ 0.09 0.35 $\pm$ 0.06 0.29 $\pm$ 0.07 0.37 $\pm$ 0.08 0.16 $\pm$ 0.08 0.49 $\pm$ 0.06 0.19 $\pm$ 0.09 0.47 $\pm$ 0.06 0.23 $\pm$ 0.08 0.48 $\pm$ 0.07
@@ 5.4 Ablation Study
6 Conclusion, Limitations, and Future Works
@@ A.2 Metrics
Algorithm 1 Training clean LDM on adversarial perturbed data with systematic defending (Summary of Sec 4.2)
1:Corrupted training set $X^{\prime}_{0}$, LDM $\theta_{0}$, CodeFormer $\phi=\{\mathcal{E}_{\phi},\mathcal{D}_{\phi},\mathcal{T}_{\phi},\mathcal{C}\}$, SR model $\psi$, Prior data $\bar{X}_{0}$, Negative Token $\mathcal{V*}_{N}$, Instance prompt $c^{ID}$, Class prompt $c$, #Generation $N_{gen}$, Quality negative prompt $\bar{\mathcal{P}_{Q}}$, Personalized identifier $\mathcal{V*}$
2:Personalized diffusion model with clean-level generation performances $\theta_{T}$, and clean personalized generation $X_{gen}$
3:// Step 1: Input Purification
4:CodeFormer: Low-quality latent $Z_{l}=\mathcal{E}_{\phi}(X^{\prime}_{0})$Rightarrow Code prediction $\tilde{Z}_{c}=\mathcal{T}(Z_{l},\mathcal{C};\phi)$Rightarrow High-quality restoration $\tilde{X}_{0}=\mathcal{D}_{\phi}(\tilde{Z}_{c})$
5:Super-resolution: Resize $\tilde{X}_{0}$ to $128^{2}$Rightarrow Applying SR model $\psi$ with quality-guided prompt and obtain $\tilde{X}_{0}^{\text{purified}}$ in $512^{2}$
6:// Step 2: Training Personalized LDM with Contrastive Decouple Learning
7:for $i$ in $1,\cdots,T$ training steps do
8:     Sample an instance data $x_{i}$ from $\tilde{X}_{0}^{\text{purified}}$, and an class-prior data $\bar{x}_{0}$ from $\bar{X}_{0}$
9:     Crafting decoupled instance prompt $c^{ID}_{dec}=\texttt{concat}(c^{ID},\mathcal{V*}_{N})$ and class-prior prompt $c_{dec}=\texttt{concat}(\mathbf{c},\text{``without''},\mathcal{V*}_{N})$.
10:     Optimize the LDM $\theta_{i}$ with standard DreamBooth loss $\mathcal{L}_{db}$triangleright Following Eq. 2
11:     $\mathcal{L}_{db}(x_{i},c^{ID}_{dec},\bar{x}_{0},c_{dec};\theta_{i})=\mathcal{L}_{\mathrm{denoise}}\left(x_{i},c^{ID}_{dec}\right)+\lambda\mathcal{L}_{\mathrm{denoise}}\left(\bar{x}_{0},c_{dec}\right),$
12:     Update LDM weight $\theta_{i}$ with $\nabla_{\theta_{i}}\mathcal{L}_{db}$ using AdamW optimizer on UNet Denoiser and Text Encoder
13:end for
14:// Step 3: Noise-free Sampling Process during Inference
15:for $j$ in $1,\cdots,N_{gen}$ do $\triangleright$ Following Eq. 6 with Classifier-free Guidance
16:     Sampling $X_{gen}^{j}$ with personalized identifier $\mathcal{V*}$, and negative prompt of $\text{concat}(\mathcal{V*}_{N},\bar{\mathcal{P}_{Q}})$ with personalized LDM $\theta_{T}$
17:end for
18:return $\theta_{T}$, $X_{gen}$
@@ Appendix B More Experimental Results
Figure 9: More visualization of perturbed and purified images.
@@ C.2 Protective Perturbation Methods
We test a wide range of protective perturbation approaches, including those that craft noise against fixed LDMs by exploiting the out-of-distribution adversarial vulnerability of DNNs [7, 57, 28, 5, 6], and those that jointly and alternatively learn the noise generator and perturbation [4, 8, 29], which show better protection capacity for the LDM fine-tuning settings [22, 23].
@@
Ensemble-based ASPL (EASPL). Since the model trainer’s pre-trained text-to-image generator is often unknown, an improved approach is to use an ensemble [61, 62] of surrogate models finetuned from different pre-trained generators, which can lead to better transferability. We implement this approach with three surrogates. Besides, we follow the practice of a single model at a time in an interleaving manner to produce optimal perturbed data due to GPU memory constraints.
MetaCloak. Despite the effectiveness of perturbation crafted from noise-surrogate joint learning, studies find that these approaches lack robustness against simple data transformations such as minor Gaussian filtering. To address this issue, MetaCloak [8] solves the underlying bi-level poisoning problem using a meta-learning framework with an additional transformation sampling process to craft transferable and robust perturbations. Incorporating an additional transformation process and a denoising-error maximization loss brings severe performance degradation in generation.